### v1

# Counterfactual Simulatability of LLM Explanations for Generation Tasks

Abstract
LLMs can be unpredictable, as even slight alterations to the prompt can cause the output to change in unexpected ways. Thus, the ability of models to accurately explain their behavior is critical, especially in high-stakes settings. One approach for evaluating explanations is counterfactual simulatability, how well an explanation allows users to infer the model’s output on related counterfactuals. Counterfactual simulatability has been previously studied for yes/no question answering tasks. We provide a general framework for extending this method to generation tasks, using news summarization and medical suggestion as example use cases. We find that while LLM explanations do enable users to better predict LLM outputs on counterfactuals in the summarization setting, there is significant room for improvement for medical suggestion. Furthermore, our results suggest that the evaluation for counterfactual simulatability may be more appropriate for skill-based tasks as opposed to knowledge-based tasks.
1 Introduction
While large language models (LLMs) have proven effective for a diverse range of applications, their outputs still often contain hallucinations of unsupported information (Ji et al., 2023) or biases (Sheng et al., 2021) that hinder their reliability on critical language generation tasks. At the same time, it is important that users themselves can accurately evaluate the model capabilities, knowledge, and reliability (Steyvers et al., 2025). Particularly in high stakes domains, such as medical applications, misunderstandings or the lack of ability to predict model behavior on unseen inputs can pose disastrous risks to users (Michalowski et al., 2024).
To anticipate these risks, recent work has turned attention to evaluating the reliability of LLM explanations (Madsen et al., 2024; Turpin et al., 2023; Kunz & Kuhlmann, 2024). In particular, Chen et al. (2024) evaluates the counterfactual simulatability of LLM explanations in yes/no question answering settings. Counterfactual simulatability is a measure of how well a model’s explanation allows humans to correctly infer the model’s predictions on simulatable counterfactuals (unseen inputs where the explanation enables the user to confidently guess the model’s output). Furthermore, according to Chen et al. (2024), the counterfactual simulatability of an explanation can be decomposed into simulation generality, a measure of the diversity of simulatable counterfactuals and simulation precision, a measure of the proportion of these counterfactuals for which humans correctly infer the model’s output. Ideal model explanations should balance both generality and precision.
Figure 1: Our evaluation pipeline. Given a model’s explanation, an LLM is prompted to generate relevant counterfactuals (right) and decompose the explanation into atomic units (left). For each unit, a human annotator verifies whether the element appears in the counterfactual (simulatability) and the counterfactual output (precision).
Counterfactual simulatability, however, is equally critical to generation tasks, where the larger space of possible outputs makes understanding a model’s decision process more challenging. To fill this gap, we formalize a framework for evaluating counterfactual simulatability in language generation tasks (figure 1). We first use LLMs to produce counterfactuals and decompose explanations into atomic units. Then, human annotators determine the simulatability and precision of each unit of a models’ explanations. Rather than evaluating the factual correctness of explanations, our framework measures the ability of LLMs to accurately describe their behavior in a generative setting. The framework evaluates whether LLMs’ explanations lead to reliable human mental models consistent with their outputs.
We apply the framework to two tasks: news summarization using CNN/DM (Nallapati et al., 2016) and medical suggestion generation using the Taiwan e-Hospital Dataset (Chen et al., 2022). These specific tasks involve different trade-offs between generality and precision. Explanations for news summarization can be highly general as models may employ similar approaches towards news documents that contain common elements (e.g., dates, quotes) while differing in content. In contrast, medical suggestion explanations can be less general as the model’s suggestion may be highly dependent on the details described by the user (i.e., slight differences in symptoms can lead to completely different suggestions). This leads to a more limited set of counterfactuals the explanations apply to, though potentially being more precise in its outputs.
With these generation tasks, we conduct an initial evaluation of LLMs’ explanations and assess where models’ explanations in these generation tasks fail.
Our contributions are summarized as follow:
  1. 1.
    We propose a framework to measure the ability of LLMs to accurately describe their behavior in a generative setting by evaluating whether LLMs’ explanations lead to mental models consistent with their outputs (counterfactual simulatability).
  2. 2.
    We assess the feasibility of using LLMs to automate human annotation in our evaluation pipeline and find that an LLM achieves significant agreement with humans.
  3. 3.
    Using our framework, we evaluate Chain-of-thought and Post-hoc explanations for multiple LLMs on two complementary tasks: news summarization and medical suggestion generation. We show that LLMs generate explanations that lead to reliable mental models in news summarization, but not medical suggestion.
2 Related work
Human mental models. Humans form mental models of the physical world as a whole (Gentner & Stevens, 2014) as well as specific technologies (e.g., Payne (1991); Du et al. (2018); Lei et al. (2016)) through their past experiences and observations. Specifically with regard to artificial intelligence systems, explanations of model predictions have been considered high quality if they provide users with an accurate and generalizable understanding of the system in the form of mental models (Rutjes et al., 2019; Merry et al., 2021). When such explanations are successful, they can improve users’ ability to effectively use AI models (Vasconcelos et al., 2022; Senoner et al., 2024) as well as to anticipate and correct undesirable model behaviors, such as biases and incorrect predictions (Bansal et al., 2019). We specifically evaluate LLMs’ explanations in generation tasks considering their ability to help form mental models.
Explanation evaluation. A variety of different approaches have been used to evaluate the quality and utility of model-generated natural language explanations and rationales. Studies evaluating natural language explanations similar to word attributions (Huang et al., 2023; Madsen et al., 2024) as well as unconstrained explanations (Turpin et al., 2023) have focused on faithfulness measures. In particular, work has proposed metrics for dimensions including comprehensiveness (DeYoung et al., 2020), sufficiency (DeYoung et al., 2020), alignment with human rationales (Fayyaz et al., 2024), and scrutability (Xu et al., 2023) among others. In contrast to intrinsic measures of model explanations, other work evaluates explanations extrinsically through impact on task performance (e.g., Camburu et al. (2018); Wei et al. (2022); Krishna et al. (2023)). Among these topics, relatively few works have investigated natural language explanations in language generation tasks; such studies include dialogue responses (Zhou et al., 2021), dialogue understanding (Gao et al., 2024), and more prominently, open-ended question answering (Ho et al., 2023; Fragkathoulas & Chlapanis, 2024; Lyu et al., 2023). While Chen et al. (2024) introduces counterfactual simulatability in classification settings, we propose a novel framework for counterfactual simulatability in language generation settings.
3 Counterfactual simulatability for generation tasks
Explanations can be evaluated by considering to what extent an observer, having seen a model’s explanation for some input, can infer (i.e., simulate) the model’s output for a counterfactual input. For instance, if a user asks an LLM “do dolphins swim?” and a model answers “yes” with the explanation “all aquatic animals swim”, then the user would infer that when asked the counterfactual “do starfish swim?”, the model will similarly answer “yes”. If, in reality, the model answers “no” to this question, then, as Chen et al. (2024) notes, the explanation is ineffective because it creates a mental model that is inconsistent with the model’s behavior despite the answer being factually correct. Counterfactual simulatability measures how accurate these mental models are on simulatable counterfactuals, unseen inputs where the explanation should allow the user to confidently guess the model’s output (Chen et al., 2024).
However, in contrast to this classification example, generation tasks have a much larger output space (e.g., there exists many possible summaries for a news document or free text responses to an open-ended medical question). This makes simulation extremely challenging as it is nearly impossible to precisely identify a single possible output based on its explanation. For instance in news summarization, if a user is shown the explanation “the summary should include the key event” and a counterfactual document on “the opening ceremony of the 2024 Paris Olympics”, while the user can logically infer that the summary will include the opening ceremony, it is impossible to predict the exact wording in which it will appear. Explanations typically cannot enable humans to pinpoint a single model output. However, they are still very useful if they help humans narrow down the possible outputs (e.g., refining “all possible summaries” to “summaries that mention the opening ceremony”).
3.1 Notation
For a given generation task, a model $M$ takes an input $x\in X$ and produces an output $o_{x}\in O$ and a corresponding explanation $e_{x}$. Here, the input, output, and explanation are all natural language and, in the case of generation, $|O|$ may be arbitrarily large. A human observes $x,e_{x}$, and forms a one-to-many mental model $h_{x,e_{x}}:X\rightarrow\mathcal{P}(O)$, where $\mathcal{P}(O)$ denotes the power set of $O$, and $h_{x,e_{x}}(x^{\prime})$ denotes what the human infers to be $M$’s possible outputs on a counterfactual $x^{\prime}$. For simplicity, $h_{e_{x}}(x^{\prime})$ is used to denote $h_{x,e_{x}}(x^{\prime})$.
3.2 Simulatability
Given a mental model, a counterfactual input is deemed simulatable if the observer can refine their expectation of the model’s output on the counterfactual (i.e., $|h_{x,e_{x}}(x^{\prime})|\ll O$). Our approach to determining whether a counterfactual is simulatable is outlined further in section 4.3.
3.3 Simulation generality and precision
The generality of an explanation is defined as the diversity of simulatable counterfactuals it applies to. In the previous example, the explanation “the summary should include the key event” is more general compared to “the summary should mention the opening ceremony of the 2024 Paris Olympics” since the former can be applied to documents covering a broader range of topics. Generality is calculated as one minus the average pairwise similarity between simulatable counterfactuals:
$\text{generality}=1-\mathbb{E}_{x^{\prime},x^{\prime\prime}\sim p,x^{\prime}\neq x^{\prime\prime}}[\alpha(x^{\prime},x^{\prime\prime})],$ (1)
where $p$ is the distribution of simulatable counterfactuals and $\alpha$ is a similarity metric such as cosine similarity. Precision measures whether the model’s actual output $M(x^{\prime})$ is within the observer’s inferred output space:
$\text{precision}=\mathbb{E}_{x^{\prime}\sim p}\left[\mathbf{1}[M({x^{\prime}})\in h_{e_{x}}(x^{\prime})]\right].$ (2)
In practice, it is unfeasible for humans to enumerate all valid outputs in $h_{e_{x}}(x^{\prime})$ due to the large output space and so determining whether the model’s output is a member of the human’s inferred output space is difficult. In section 4.3, we propose a method to estimate mental models by evaluating the presence of inferred atomic units of information in the model’s output.
4 Methods
Figure 2: Example explanations, counterfactuals, counterfactual outputs, and annotations for news summarization and medical suggestion. Atomic units of the explanation are highlighted (for medical suggestion, blue: patient information, orange: suggestions).
4.1 Datasets
We select two generative tasks that represent practical LLM use cases and whose natural language explanations differ in relation to generality and precision. For the first task of news summarization, we use the CNN/DM dataset (Nallapati et al., 2016). Summarization is chosen as it is a well-studied application for language generation. Additionally, this task allows us to study high-level, abstract explanations since summarization does not rely on facts learned during the LLM’s pre-training, rather its ability to identify key elements from the input document.
Conversely, we also investigate medical suggestion, a domain where faithful explanations are critical as they allow humans to accurately predict the model’s behavior. For this task, we use the Taiwan e-Hospital Dataset, a collection of 86,399 Mandarin question answer pairs from the online health website Taiwan e-Hospital (Chen et al., 2022). We select this specific dataset because each sample consists of a question, suggestion, explanation triple, ensuring that the questions are complex enough such that explanations are helpful. Furthermore, this dataset mirrors the practical setting we aimed to study, where everyday users interact with LLMs to seek general medical advice without any specialized knowledge.
The differences between explanations for these tasks are illustrated in figure 2. In contrast to summarization, generating medical suggestions is more knowledge-based, requiring the LLM to identify key elements of the user’s query (e.g. their expressed symptoms, any relevant medical history) and relate them to potential suggestions using knowledge encoded in the model. As such, the explanations are highly specific to the input. For instance in figure 2, whereas the summarization explanation applies to counterfactuals where abstract elements such as a “key event” or “quotes or states’ are mentioned, the medical suggestion explanation can only be used to infer the model’s output when a user expresses “persistent nasal discharge”, “frequent sneezing”, and “symptoms [that] have lasted for over three months”. While explanations for medical suggestion are likely to apply to a more limited range of counterfactuals, they allow the user to infer specific pieces of information in the output as opposed to high-level elements as is the case for summarization. Therefore, these two tasks are complementary for studying counterfactual simulatability as they represent different requirements for generality and precision.
4.2 Explanations
Through prompting (instructions, few-shot examples), we guide the LLMs to generate explanations that emphasize this difference in generality and precision across tasks. For news summarization, to be highly generalizable, we encourage the model to explain its decision process at a high-level, identifying abstract elements while avoiding reference to specific topics. For medical suggestion, we instruct the model to first identify important details in the user’s question, propose possible underlying causes, and suggest recommended actions accordingly. These approaches reflect the skill-based vs. knowledge-based nature of our tasks. Additionally, following Chen et al. (2024), we experiment with both Chain-of-thought and Post-hoc prompting to produce explanations (Camburu et al., 2018). The prompts used are provided in appendix F.
4.3 Estimating mental models
Generation tasks are particularly challenging for counterfactual simulatability because mental models are extremely difficult to estimate given the large space of possible outputs. To this end, we introduce a method for estimating mental models by considering atomic units of information in the explanation as a proxy.
Intuitively, explanations $e_{x}$ identify key pieces of information that the model considers essential for the output. Thus, an observer may simulate that if the counterfactual input $x^{\prime}$ includes a piece of information $a$ deemed important by the model in $e_{x}$, the model should similarly take into account $a$ in its output for $x^{\prime}$. Based on this explanation $e_{x}$, an observer’s mental model $h_{e_{x}}$ can be formalized as:
$h_{e_{x}}(x^{\prime})=\{o\in O\mid\forall a\in e_{x}\cap x^{\prime},a\in o\},$ (3)
where the space of possible outputs on the counterfactual is refined to only outputs that contain the elements $\{a\}$. With this, we formulate simulatability and precision as:
$\text{simulatability}=\mathbb{1}\{\forall a\in e_{x},a\in x^{\prime}\}$ (4)
$\text{precision}=\frac{|\{a\in e_{x}\cap M(x^{\prime})\}|}{|\{a\in e_{x}\}|}$ (5)
First, we determine whether a counterfactual is simulatable by verifying whether all atomic units of the explanation appear in the counterfactual (equation 4). Then, for simulatable counterfactuals, precision is measured as the proportion of these atomic units that appear in the model’s counterfactual output (equation 5). Generality is calculated using the approach outlined in section 3.3 using cosine similarity. Finally, we average generality and precision scores across explanations.
A key distinction between the tasks is that while each atomic unit of a summarization explanation is linked to an expected item in the input (for simulatability) and its reference in the output (for precision), medical suggestion explanations do not follow this one-to-one mapping of units between input and output. For instance, using the example provided in figure 2, while ”Persistent nasal discharge” is identified as important, it is unclear which suggested action the model has generated in response to this symptom. To address this, for the medical domain, we further instruct the LLM when parsing the explanation to classify units into two categories, patient information and suggestions. These categories are then used to evaluate simulatability and precision respectively during annotation (see figure 2).
4.4 Experimental setup
Human evaluation.
Figure 1 provides a high-level overview of our evaluation framework. First, using randomly sampled inputs from our dataset, we prompt GPT-4 Turbo (see section 4.2), to generate outputs and explanations. For each of these explanations, we use GPT-4 Turbo to generate 3 relevant counterfactuals, providing the explanation in the prompt as context. Then, we use GPT-4 Turbo to decompose the explanation into atomic units (i.e., $e_{x}\rightarrow\{a\}$) and, in the medical suggestion use case, instruct it to extract units into two groups for patient details or suggested actions. We opted for GPT-4 Turbo as the LLM in our evaluation framework due to its performance and relatively low cost. Finally, we generate counterfactual outputs using the same prompt as the first step and have humans examine the counterfactual and counterfactual output to annotate simulatability and precision for each atomic unit of the explanation. To evaluate the ability of GPT-4 Turbo to break down explanations, we also instruct the annotators to indicate whether each unit was extracted correctly and whether any details that should have been extracted are missing. Furthermore, we investigate if increasing the number of counterfactuals generated per explanation affects generality, but find no noticeable differences (appendix E). Additional details for the human evaluation and the prompts used are provided in appendix A and F respectively.
Comparison of human-LLM annotations.
We evaluate whether using an LLM to automate human annotation is feasible by prompting GPT-4 Turbo with similar instructions and assessing the inter-annotator agreements between human-human and human-LLM pairs.
Automatic evaluation.
Finally, using GPT-4 Turbo for annotation, we repeat our framework and conduct a larger evaluation of three popular LLMs using more data and five counterfactuals per explanation. Additional details for the automatic evaluation are provided in appendix B.
5 Results
5.1 Intermediate results
GPT-4 Turbo is able to parse explanations well for summarization but not for medical suggestion.
Annotators find that the LLM successfully breaks down explanations with 96% and 57% accuracy for news summarization and medical suggestion respectively. For medical suggestion errors, the LLM either fails to identify a key detail or extracts an atomic unit incorrectly (extracted a unit it should not have or classified it as the incorrect type of information). A detailed break down of parsing errors is provided in appendix C.
GPT-4 Turbo is able to generate simulatable counterfactuals for summarization but not for medical suggestion.
While almost all ($74/76$) generated counterfactuals are deemed simulatable in the summarization setting, only slightly more than half are for medical suggestion ($52/90$). Most non-simulatable counterfactuals for medical suggestion contain $0.4-0.8$ of the atomic units of the explanation, indicating that counterfactual generation is more difficult in this setting. The distribution of generated counterfactuals, bucketed by the proportion of atomic units that appear is provided in appendix D.
5.2 Human evaluation
Task Explanation Method Generality Precision
News Summarization Chain of Thought 0.52 0.81
Post Hoc 0.49 0.89
Medical Suggestion Chain of Thought 0.20 0.51
Post Hoc 0.26 0.59
Table 1: Generality and precision across explanation types and tasks for human evaluation.
GPT-4 Turbo explanations lead to generalizable counterfactuals and consistent mental models in the case of summarization, with approximately $0.8$ of the inferred information appearing in the counterfactual output on average. In contrast, explanations in the medical suggestion setting are less generalizable and less precise (approximately $0.5$), indicating that LLMs may struggle more to reliably explain their behavior for this task. Chain-of-thought and Post-hoc explanations lead to similar results (see table 1).
Additionally, as a sanity check, we generate counterfactual outputs conditioned on the original explanation and verify that these result in a precision score of $1.00$. Since we expect that the model will perfectly follow the explanation, this precision score indicates that our framework does indeed capture how well models follow their explanations. Thus, the low precision scores for medical suggestion are due to the model’s behavior not adhering to its explanation on that task rather than the evaluation setup.
5.3 Automatic evaluation
News Summarization Medical Suggestion
Annotator 1 2 3 GPT-4 Turbo n 1 2 3 GPT-4 Turbo n
1 - 0.35 0.71 0.64 263 - 0.76 0.74 0.68 273
2 0.35 - 0.57 0.48 213 0.76 - 0.73 0.54 261
3 0.71 0.57 - 0.71 73 0.74 0.73 - 0.74 258
Table 2: Inter-annotator agreement (Cohen’s Kappa) between human annotators and GPT-4 Turbo for news summarization and medical suggestion.
GPT-4 Turbo is able to approximate human annotation for our tasks.
Table 2 reports the pairwise Cohen’s Kappa between each pair of annotators and the LLM. We find that overall, GPT-4 Turbo achieves similar inter-annotator agreement to humans, with an average of 0.61 for human-LLM pairs compared to 0.54 for human-human pairs in news summarization and 0.65 for human-LLM pairs compared to 0.74 for human-human pairs in medical suggestion.
Task  Model Chain-of-thought Post-hoc
# Expl # Samples Generality Precision # Expl # Samples Generality Precision
News Summarization Claude 3.7 Sonnet 48 189 0.67 0.93 45 173 0.62 0.84
GPT-4 46 160 0.59 0.84 48 193 0.65 0.78
Llama 3 47 172 0.67 0.74 43 153 0.67 0.66
Medical Suggestion Claude 3.7 Sonnet 24 55 0.21 0.48 24 78 0.20 0.66
GPT-4 36 103 0.20 0.46 36 122 0.19 0.65
Llama 3 24 69 0.19 0.56 30 110 0.20 0.66
Table 3: Generality and precision across models, explanation types, and tasks for automatic evaluation.
The results of the automatic evaluation, presented in table 3, further support our findings from the human evaluation. Namely, models are much better able to accurately explain their behavior for summarization as opposed to medical suggestion while remaining general such that these explanations apply to a diverse set of counterfactuals. Chain-of-thought explanations tend to lead to better mental models for summarization, which may reflect the skill-based nature of this task. On the other hand, Post-hoc explanations lead to more precise explanations for medical suggestion. Models may also differ in their ability to describe their behavior in a generative setting, for instance, Claude 3.7 Sonnet demonstrates noticeably better precision scores for news summarization compared to other models.
6 Discussion
LLM explanations are helpful for skill-based generation tasks but struggle for knowledge-based generation tasks.
While model explanations do enable users to infer pieces of information that will appear in counterfactual outputs in both settings, the mental models they produce are more accurate in summarization compared to medical suggestion. Furthermore, summarization explanations, employing a high-level approach in their explanation, also lead to counterfactuals that are more generalizable. These may suggest that LLMs are better able to describe their behavior for skill-based tasks, where users can reliably infer the elements that will appear in the outputs, compared to knowledge-based tasks, where users are less capable of reliably inferring specific points (e.g., suggestions). Additionally, using Chain-of-thought prompting, which aligns with the skill-based nature of summarization, may lead to more precise explanations.
These findings may reflect the predictability of model behavior on different tasks types. For medical question, minor alterations to the question may lead to very different answers. For instance, if a user describes ”I experience chest pain sometimes when I exercise.” a model might respond ”consider reducing exercise intensity, try warming up thoroughly before exercise …”. However, if a user changes their question slightly to ”I experience chest pain sometimes when I exercise. I started taking a new pre-workout supplement with high caffeine recently.”, although they are expressing the same symptoms, because of the presence of additional information the model might respond ”discontinue the supplement or use smaller dosages, the chest pain is likely related to caffeine-induced heart palpitations …”. These differences in the sensitivity of the model’s output relative to the input may lead to the differences in precision and generality observed in our experiments.
Our counterfactual simulatability evaluation framework is effective in the summarization setting but less suited for medical suggestion.
Although LLMs are able to automate the human annotation steps in our evaluation, they demonstrate issues in other aspects for the medical suggestion task. Specifically, annotators identify many errors in the parsing of explanations into atomic units, where the LLM misses key information or mis-classifies the information. In one instance, “possible heat-induced asthma” is incorrectly extracted as an expressed symptom when in reality it is a potential cause. Unlike the summarization setting, this additional requirement of classifying the type of information introduces more complexity, making the LLM a less effective tool in the evaluation pipeline. Additionally, we found that the LLM is unable to produce as many simulatable counterfactuals in the medical domain compared to summarization. There is significant room for improvement towards adapting counterfactual simulatability for knowledge-based tasks like medical suggestion, where the explanation and decision process for an LLM is less explicit and relies heavily on knowledge encoded in the model.
Appendix A Human Evaluation
Annotators: Two groups of 3 student annotators each from the Columbia NLP lab were recruited for the human evaluation (one group per task). Since these tasks were meant to replicate LLMs in a practical setting, where a general user is asking for a brief summary of an article or seeking some quick medical advice, the annotators were not required to have any specialized domain knowledge.
Distribution of Annotations: For each task and explanation type (Chain-of-thought and Post-hoc), annotators were asked to annotate across a total of 15 explanations with 3 counterfactuals each, resulting in 45 samples (explanation, counterfactual pairs). In order to compare human-human agreement against human-LLM agreement, each annotator was assigned the same set of 3 samples plus an additional 4. Not all annotations were completed for news summarization, leading to a slight discrepancy in annotations across tasks.
News Summarization Instructions:
  • Parsed Explanation Annotations: Check if the explanation (Column B) is parsed correctly in Column C. If yes, put ”ok”. If not, write which information in the explanation is missing / which info in column C are hallucinating.
  • Document Annotations: Check whether each item in the parsed explanation appears in the document or not. Put either ”no” or ”yes”.
  • Summary Annotations: Check whether each item in the parsed explanation appears in the summary or not. Put either ”no” or ”yes”
Medical Suggestion Instructions:
Tasks 1, 3:
  • for each of the ”Extracted Points”, label ”Y” or ”N” if the point was parsed correctly from the AI explanation
  • if there are missing points or other comments, feel free to note them in the EXTRA QUESTION annotation cell
  • task 1 focuses on extracting patient details (e.g., symptoms, medical history)
  • task 3 focuses on extracting suggested next steps (e.g., recommended treatments)
  • since breakdowns are the same across rows, feel free to complete just one per example ID
Tasks 2, 4:
  • for each of the ”Extracted Points”, label ”Y” or ”N” if the point is reflected in the counterfactual (task 2) or AI suggestion (task 4)
  • the matching is flexible, i.e., point does not necessarily need to appear in exactly the same wording to be marked ”Y”
  • for example, in example 3 the point mentions ””hemoglobin level of 10.5, slightly below normal”” these specific scores do not need to be mentioned, rather ””below normal levels”” is sufficient.
Appendix B Automatic Evaluation
The automatic evaluation consisted of 50 explanations per task and generating 5 counterfactuals each (250 explanation, counterfactual pairs total). The models evaluated were: Anthropic’s Claude 3.7 Sonnet, OpenAI’s GPT-4, and Meta’s Llama 3.3 70B Instruct. These models were chosen as they represent a few popular LLM options as well as a mix of proprietary and open-source models.
Appendix C Explanation Parsing Quality Results
We show a categorized break down of the reported errors for explanation parsing in both tasks.
Parsed Explanations News Summarization Medical Suggestion
($n=26$) ($n=30$)
Accuracy 0.96 0.57
Breakdown of Incorrect Examples:
      Missing Extraction 1 8
      Incorrect Extraction 0 4
      Missing and Incorrect Extraction 0 1
Table 4: Observed errors in explanation parsing across tasks.
Appendix D Counterfactual Generation: Simulatability
We calculate the proportion of atomic units that appear in each counterfactual for the human evaluation (where 1.00 indicates a simulatable counterfactual), bucket them, and display the results below.
Proportion of atomic units present News Medical
in counterfactual (human annotated) Summarization Suggestion
1.00* 74 52
0.80–0.99 2 0
0.60–0.79 2 20
0.40–0.59 0 17
0.20–0.39 0 1
0.00–0.19 0 0
Total 76 90
Table 5: Distribution of the proportion of atomic units present in GPT4-Turbo generated counterfactuals across tasks. *Indicates the set of simulatable counterfactuals.
Appendix E Counterfactual Generation: Generality
We show how generality and the number of simulatable counterfactuals changes as we adjust the number of counterfactuals generated per explanation. For news summarization and 10 counterfactuals per explanation, not all 100 counterfactuals were generated due to errors parsing the LLMs output. Since news summarizations are more expensive to produce (longer context prompt), we only use 10 explanations compared to 30 for medical suggestion.
Metric Counterfactuals generated per explanation
3 5 10
News Summarization
Generality 0.497 0.512 0.595
Simulatable counterfactuals 18 30 38
Total generated counterfactuals 30 50 80
Medical Suggestion
Generality score 0.187 0.218 0.227
Simulatable counterfactuals 41 66 95
Total generated counterfactuals 90 150 300
Table 6: Generality and simulatability as the number of generated counterfactuals per explanation increases.
Appendix F LLM Prompts
We provide the prompts used in our evaluation pipeline.
Figure 3: News Summarization: Chain-of-thought explanations
Figure 4: News Summarization: Post-hoc explanations
Figure 5: News Summarization: Counterfactual generation
Figure 6: News Summarization: Explanation extraction
Figure 7: News Summarization: Simulatability annotation
Figure 8: News Summarization: Precision annotation
Figure 9: Medical Suggestion: Chain-of-thought explanations
Figure 10: Medical Suggestion: Post-hoc explanations
Figure 11: Medical Suggestion: Counterfactual generation
Figure 12: Medical Suggestion: Explanation extraction
Figure 13: Medical Suggestion: Simulatability annotation
Figure 14: Medical Suggestion: Precision annotation